# End-to-end Autonomous Driving: Challenges and Frontiers

### v1

@@ 4.7 Causal Confusion
Driving is a task that exhibits temporal smoothness, which makes past motion a reliable predictor of the next action. However, methods trained with multiple frames can become overly reliant on this shortcut [211] and suffer from catastrophic failure during deployment. This problem is referred to as the copycat problem [57] in some works and is a manifestation of causal confusion [212], where access to more information leads to worse performance. One of the earliest reports of this effect was made by LeCun et al. [213]. They used a single input frame for steering prediction to avoid such extrapolation. Though simplistic, this is still the preferred solution in current state-of-the-art imitation learning methods [28, 22, 29]. Unfortunately, using a single frame has a downside of not being able to extract the velocity of surrounding actors. Another source of causal confusion is speed measurement [16]. Fig. 7 showcases an example of a car waiting at a red light. The speed of the car is highly correlated with the brake action because the car is waiting for many frames where the speed is zero and the action is the brake. Only at the single frame where the traffic light changes from red to green does this correlation break down.
@@
There are multiple approaches to combat the causal confusion problem when using multiple frames. ChauffeurNet [214] addresses the issue by using intermediate visual abstractions in BEV. One abstraction is the ego agent’s past, while other abstractions do not contain this information. During training, the past motion of the ego agent is dropped out with a 50% probability. However, explicit abstractions are required for this approach to work effectively. In [57], the authors attempt to remove spurious temporal correlations from a learned intermediate bottleneck representation by training an adversarial model that predicts the ego agent’s past action. This results in a min-max optimization problem where the imitation loss is minimized, and the adversarial loss is maximized. Intuitively, this trains the network to eliminate its own past from its intermediate layers. This approach works well in MuJoCo but does not scale to complex vision-based driving. The first to work on driving complexity is [58]. They propose upweighting keyframes in the training loss. Keyframes are frames where a decision change occurs (and hence are not predictable by extrapolating the past). To find the keyframes, they train a policy that predicts the action with only the ego agent’s past as input. PrimeNet [60] improves performance compared to keyframes by using an ensemble, where the prediction of a single-frame model is given as additional input to a multi-frame model. Chuang et al. [215] do the same but supervise the multi-frame network with action residuals instead of actions. OREO [59] maps images to discrete codes representing semantic objects and applies random dropout masks to units that share the same discrete code. This helps in Confounded Atari, where the previous action is rendered on the screen. In autonomous driving, the problem of causal confusion can be circumvented by using only LiDAR histories (with a single frame image) and realigning point clouds into the same coordinate system. This removes information about ego-motion while retaining information about other vehicles’ past states. This technique has been used in multiple works [32, 1, 52], though it was not motivated this way.
@@
Causal confusion in imitation learning has been a persistent challenge for nearly two decades. Significant effort has gone into studying this problem in recent years [59, 215, 60, 58, 57, 212]. However, these studies have used environments that are modified to simplify studying the causal confusion problem. Showing performance improvements in state-of-the-art settings remains an open problem.
4.8 Robustness
@@ 4.8.1 Long-tailed Distribution
One important aspect of the long-tailed distribution problem is dataset imbalance, where a few classes make up the majority and many other classes only have a limited number of samples, as shown in Fig. 8 (a). This poses a big challenge for models to generalize to various environments. Various methods tackle this problem with data processing, including over-sampling [216, 217, 218, 219, 220], under-sampling [221, 222, 223], and data augmentation [224, 225, 226]. Besides, weighting-based approaches [227, 228, 229, 230, 231, 232, 233] are also commonly used to mitigate the dataset imbalance problem.
In the context of end-to-end autonomous driving, the issue of long-tailed distributions is particularly severe. Dataset imbalance is especially problematic in driving datasets because most typical drives are repetitive and uninteresting e.g., following a lane for many frames. Conversely, interesting safety-critical scenarios occur rarely but are diverse in nature. To tackle this issue, some works rely on handcrafted scenarios [234, 13, 235, 90, 99] to generate more diverse and interesting data in simulation. LBC [5] leverages the privileged agent to create imaginary supervisions conditioned on different navigational commands. LAV [52] argues that though the ego car for data collection has few accident-prone cases, other agents may have experienced some safety-critical or interesting cases. Therefore, it includes trajectories of other agents for training to promote data diversity. In [236], a simulation framework is proposed to apply importance-sampling strategies to accelerate of the evaluation of rare-event probabilities.
Another line of research [86, 35, 237, 34, 7] generates safety-critical scenarios in a data-driven manner through adversarial attacks. In [86], Bayesian Optimization is employed to generate adversarial scenarios. Learning to collide [35] represents driving scenarios as the joint distribution over building blocks and applies policy gradient RL methods to generate risky scenarios. AdvSim [34] modifies agents’ trajectories, while still adhering to physical plausibility, to cause failures and updates the LiDAR accordingly. The recent work KING [7] proposes an optimization algorithm for safety-critical perturbations using gradients through differentiable kinematics models.
@@ 4.8.2 Covariate Shift
As discussed in Sec. 2.1, one important challenge for behavior cloning is covariate shift. The state distributions from the expert’s policy and those from the trained agent’s policy differ, leading to compounding errors when the trained agent is deployed in the unseen testing environment or when the reactions from other agents differ from training time. This could result in the trained agent being in a state that is outside the expert’s distribution for training, leading to severe failures. An illustration is presented in Fig. 8 (b). DAgger (Dataset Aggregation) [26] is a commonly used method to overcome this issue. DAgger is an iterative training process, in which the current trained policy is rolled out in each iteration to collect new data, and the expert is used to label the visited states. This enriches the training dataset by adding examples of how to recover from the suboptimal states that an imperfect policy might visit. The policy is then trained on the augmented dataset, and the process repeats. However, one downside of DAgger is the need for an available expert to query online.
For end-to-end autonomous driving, DAgger is adopted in [24] with an MPC-based expert. To reduce the cost of constantly querying the expert and improve safety, SafeDAgger [25] extends the original DAgger algorithm by learning a safety policy that estimates the deviation between the current policy and the expert policy. The expert is only queried when the deviation is large, and the expert takes over in those dangerous cases. MetaDAgger [56] combines meta-learning with DAgger to aggregate data from multiple environments. LBC [5] adopts DAgger and resamples the data so that samples with higher loss are sampled more frequently. In DARB [10], several modifications are made on DAgger to fit the driving task. To better utilize failure or safety-related samples, it proposes several mechanisms, including task-based, policy-based, and policy & expert-based mechanisms, to sample such critical states. It also uses a fixed-sized replay buffer for iterative training to increase diversity and reduce dataset bias.
@@ 4.8.3 Domain Adaptation
  • Day-to-night: illumination variations in the sensor input.
@@
VISRI [239] uses a translation network to map simulated images to real images, with segmentation maps serving as the intermediate representation. An RL agent is trained based on the translated simulated images. In [240], domain-invariant feature learning is achieved with image translators and discriminators to map images from two domains into a common latent space. Similarly, LUSR [241] adopts a Cycle-Consistent VAE to project images into a latent representation comprised of a domain-specific part and a domain-general part, based on which the policy is learned. UAIL [209] targets weather-to-weather adaptation by disentangling the image from different weather conditions into a distinguishable style space and a shared content space with a GAN. In SESR [169], class disentangled encodings are extracted from a semantic segmentation mask to reduce the domain gap between images in simulators and the real world. Domain randomization [242, 243, 244] is also a simple and effective technique for sim-to-real adaptation for RL policy learning in simulators, which is further adapted for end-to-end autonomous driving [245, 167]. It is realized by randomizing the rendering and physical settings of the simulators to cover the variability of the real world during training, and obtain a trained policy with good generalization ability.
Currently, sim-to-real adaptation through source target image mapping or domain-invariant feature learning is the focus for end-to-end autonomous driving. Other domain adaptation cases such as geography-to-geography or weather-to-weather adaptation are handled through the diversity and scale of the training dataset. As LiDAR has become a popular input modality for driving, specific adaptation techniques tailored to the characteristics of LiDAR must also be designed, given that current works mainly concentrate on image-based adaptation. Besides, attention should be drawn to traffic agents’ behavior and traffic rule gaps between the simulator and the real world, as current methods only focus on the visual gap in images. Incorporating real-world data into simulation through techniques such as NeRF [113, 246] is another promising direction.
@@ 5.1 Zero-shot and Few-shot Learning
It is inevitable for autonomous driving models to eventually encounter real-world scenarios that lie beyond the training data distribution. This raises the question of whether we can successfully adapt the model to an unseen target domain where limited or no labeled data is available. Formalizing this task for the end-to-end driving domain and incorporating techniques from the zero-shot/few-shot learning literature are the key steps toward achieving this.
@@ 5.2 Modular End-to-end Planning
The modular end-to-end planning framework optimizes multiple modules while prioritizing the downstream planning task, which enjoys the advantages of interpretability as indicated in Sec. 4.6. This is advocated in recent literature [247, 2] and certain industry solutions (Tesla, Wayve, etc.) have involved similar ideas. When designing these differentiable perception modules, several questions arise regarding the choice of loss functions, such as the necessity of 3D bounding boxes for object detection, whether the occupancy representation is sufficient for detecting general obstacles, or the advantages of opting for BEV segmentation over lane topology for static scene perception.
@@ 5.3 Data Engine
The importance of large-scale and high-quality data for autonomous driving can never be emphasized enough. Establishing a data engine with an automatic labeling pipeline [248] could greatly facilitate the iterative development of both data and models. The data engine for autonomous driving, especially modular end-to-end planning systems, needs to streamline the process of annotating high-quality perception labels with the aid of large perception models in an automatic way. It should also support mining hard/corner cases, scene generation, and editing to facilitate the data-driven evaluations discussed in Sec. 3.1 and promote diversity of data and the generalization ability of models (Sec. 4.8). A data engine would enable autonomous driving models to make consistent improvements.
@@ 5.4 Foundation Model
Recent advancements in large foundation models in both language [164, 165, 249, 250] and vision [251, 252, 248, 253] have had a significant impact on various aspects of society. The utilization of large-scale data and model capacity has unleashed the immense potential of AI in high-level reasoning tasks. The paradigm of finetuning [254, 255] or prompt learning [163], optimization in the form of self-supervised reconstruction [256] or contrastive pairs [161], and data pipelining, etc., are all applicable to the end-to-end autonomous driving domain. However, we contend that the direct adoption of LLMs for autonomous driving might seem misaligned between the different goals of the two targets. The output of an autonomous agent often requires steady and accurate measurements, whereas the generative sequence output in language models aims to behave like a human, irrespective of its accuracy. A feasible solution to develop a large autonomous driving model is to train a video predictor that can forecast long-term predictions of the environment, either in 2D or 3D. To perform well on downstream tasks like planning, the objective to be optimized for the large model needs to be sophisticated enough, beyond frame-level perception.
5.5 Vehicle-to-everything (V2X)
Occlusion and obstacles beyond the perceptual range are two fundamental challenges for modern computer vision techniques, which can even pose great difficulties for human drivers when they need to react quickly to crossing agents. Vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-everything (V2X) systems offer promising solutions to address this crux, where information from various viewpoints complements the ego blind spots. Witnessing the progress in information transmission mechanisms for multi-agent scenarios [257, 258, 259, 260, 261], these systems could provide a solution to realize advanced decision intelligence among autonomous vehicles.
6 Conclusion
In this survey, we give an overview of fundamental methodologies and summarize various aspects of simulation and benchmarking. We thoroughly analyze the extensive literature to date, and highlight a wide range of critical challenges and promising resolutions. Future endeavors to embrace the rapidly developed foundation models and data engines are discussed in the end. End-to-end autonomous driving faces great opportunities and challenges simultaneously, with the ultimate goal of building generalist agents. In this era of emerging technologies, we hope that this survey would serve as a starting point to shed new light on this domain.
@@ Acknowledgments
Li Chen, Penghao Wu, and Hongyang Li were partially supported by National Key R&D Program of China (2022ZD0160100), Shanghai Committee of Science and Technology (21DZ1100100), and NSFC (62206172). Andreas Geiger and Bernhard Jaeger were supported by the ERC Starting Grant LEGO-3D (850533), the BMWi in the project KI Delta Learning (project number 19A19013O) and the DFG EXC number 2064/1 - project number 390727645. Kashyap Chitta was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Bernhard Jaeger and Kashyap Chitta.

### v2

@@ 4.8 Causal Confusion
Driving is a task that exhibits temporal smoothness, which makes past motion a reliable predictor of the next action. However, methods trained with multiple frames can become overly reliant on this shortcut [241] and suffer from catastrophic failure during deployment. This problem is referred to as the copycat problem [57] in some works and is a manifestation of causal confusion [242], where access to more information leads to worse performance.
Causal confusion in imitation learning has been a persistent challenge for nearly two decades. One of the earliest reports of this effect was made by LeCun et al. [243]. They used a single input frame for steering prediction to avoid such extrapolation. Though simplistic, this is still a preferred solution in current state-of-the-art IL methods [28, 22]. Unfortunately, using a single frame makes it hard to extract the motion of surrounding actors. Another source of causal confusion is speed measurement [16]. Fig. 7 showcases an example of a car waiting at a red light. The action of the car could highly correlate with its speed because it has waited for many frames where the speed is zero and the action is the brake. Only when the traffic light changes from red to green does this correlation break down.
@@
There are several approaches to combat the causal confusion problem when using multiple frames. In [57], the authors attempt to remove spurious temporal correlations from the bottleneck representation by training an adversarial model that predicts the ego agent’s past action. Intuitively, the resulting min-max optimization trains the network to eliminate its past from intermediate layers. It works well in MuJoCo but does not scale to complex vision-based driving. OREO [59] maps images to discrete codes representing semantic objects and applies random dropout masks to units that share the same discrete code, which helps in confounded Atari. In end-to-end driving, ChauffeurNet [244] addresses the causal confusion issue by using the past ego-motion as intermediate BEV abstractions and dropping out it with a 50% probability during training. Wen et al. [58] propose upweighting keyframes in the training loss, where a decision change occurs (and hence are not predictable by extrapolating the past). PrimeNet [60] improves performance compared to keyframes by using an ensemble, where the prediction of a single-frame model is given as additional input to a multi-frame model. Chuang et al. [245] do the same but supervise the multi-frame network with action residuals instead of actions. In addition, the problem of causal confusion can be circumvented by using only LiDAR histories (with a single frame image) and realigning point clouds into one coordinate system. This removes ego-motion while retaining information about other vehicles’ past states. This technique has been used in multiple works [32, 1, 52], though it was not motivated this way.
@@
However, these studies have used environments that are modified to simplify studying the causal confusion problem. Showing performance improvements in state-of-the-art settings as mentioned in Sec. 3.2.5 remains an open problem.
4.9 Lack of Robustness
@@ 4.9.1 Long-tailed Distribution
One important aspect of the long-tailed distribution problem is dataset imbalance, where a few classes make up the majority, as shown in Fig. 8 (a). This poses a big challenge for models to generalize to diverse environments. Various methods mitigate this issue with data processing, including over-sampling [246, 247], under-sampling [248, 249], and data augmentation [250, 251]. Besides, weighting-based approaches [252, 253] are also commonly used.
In the context of end-to-end autonomous driving, the long-tailed distribution issue is particularly severe. Most drives are repetitive and uninteresting e.g., following a lane for many frames. Conversely, interesting safety-critical scenarios occur rarely but are diverse in nature, and hard to replicate in the real world for safety reasons. To tackle this, some works rely on handcrafted scenarios [254, 13, 255, 256, 100] to generate more diverse data in simulation. LBC [5] leverages the privileged agent to create imaginary supervisions conditioned on different navigational commands. LAV [52] includes trajectories of non-ego agents for training to promote data diversity. In [257], a simulation framework is proposed to apply importance-sampling strategies to accelerate the evaluation of rare-event probabilities.
Another line of research [258, 35, 259, 34, 7, 260] generates safety-critical scenarios in a data-driven manner through adversarial attacks. In [258], Bayesian Optimization is employed to generate adversarial scenarios. Learning to collide [35] represents driving scenarios as the joint distribution over building blocks and applies policy gradient RL methods to generate risky scenarios. AdvSim [34] modifies agents’ trajectories to cause failures, while still adhering to physical plausibility. KING [7] proposes an optimization algorithm for safety-critical perturbations using gradients through differentiable kinematics models.
@@ 4.9.2 Covariate Shift
As discussed in Sec. 2.1, one important challenge for behavior cloning is covariate shift. The state distributions from the expert’s policy and those from the trained agent’s policy differ, leading to compounding errors when the trained agent is deployed in unseen testing environments or when the reactions from other agents differ from training time. This could result in the trained agent being in a state that is outside the expert’s distribution for training, leading to severe failures. An illustration is presented in Fig. 8 (b).
DAgger (Dataset Aggregation) [26] is a common solution for this issue. DAgger is an iterative training process. The current trained policy is rolled out in each iteration to collect new data, and the expert is used to label the visited states. This enriches the dataset by adding examples of how to recover from suboptimal states that an imperfect policy might visit. The policy is then trained on the augmented dataset, and the process repeats. However, one downside of DAgger is the need for an available expert to query online.
For end-to-end autonomous driving, DAgger is adopted in [24] with an MPC-based expert. To reduce the cost of constantly querying the expert, SafeDAgger [25] extends the original DAgger algorithm by learning a safety policy that estimates the deviation between the current policy and the expert policy. The expert is only queried when the deviation is large. MetaDAgger [56] uses meta-learning with DAgger to aggregate data from multiple environments. LBC [5] adopts DAgger and resamples the data with higher loss more frequently. In DARB [10], to better utilize failure or safety-related samples, it proposes several mechanisms, including task-based, policy-based, and policy & expert-based mechanisms, to sample such critical states.
@@ 4.9.3 Domain Adaptation
  • Day-to-night: illumination variations in visual inputs.
@@
Typically, domain-invariant feature learning is achieved with image translators and discriminators to map images from two domains into a common latent space or representations like segmentation maps [262, 263]. LUSR [264] and UAIL [235] adopt a Cycle-Consistent VAE and GAN, respectively, to project images into a latent representation comprised of a domain-specific part and a domain-general part. In SESR [190], class disentangled encodings are extracted from a semantic segmentation mask to reduce the sim-to-real gap. Domain randomization [265, 266, 267] is also a simple and effective sim-to-real technique for RL policy learning, which is further adapted for end-to-end autonomous driving [268, 188]. It is realized by randomizing the rendering and physical settings of the simulators to cover the variability of the real world during training.
Currently, sim-to-real adaptation through source target image mapping or domain-invariant feature learning is the focus. Other DA cases are handled by constructing a diverse and large-scale dataset. Given that current methods mainly concentrate on the visual gap in images, and LiDAR has become a popular input modality for driving, specific adaptation techniques tailored for LiDARs must also be designed. Besides, traffic agents’ behavior gaps between the simulator and the real world should be noticed as well. Incorporating real-world data into simulation through techniques such as NeRF [113] is another promising direction.
@@ 5.1 Zero-shot and Few-shot Learning
It is inevitable for autonomous driving models to eventually encounter real-world scenarios that lie beyond the training data distribution. This raises the question of whether we can successfully adapt the model to an unseen target domain where limited or no labeled data is available. Formalizing this task for the end-to-end driving domain and incorporating techniques from the zero-shot/few-shot learning literature are the key steps toward achieving this [269, 270].
@@ 5.2 Modular End-to-end Planning
The modular end-to-end planning framework optimizes multiple modules while prioritizing the ultimate planning task, which enjoys the advantages of interpretability as indicated in Sec. 4.6. This is advocated in recent literature [271, 2] and certain industry solutions (Tesla, Wayve, etc.) have involved similar ideas. When designing these differentiable perception modules, several questions arise regarding the choice of loss functions, such as the necessity of 3D bounding boxes for object detection, whether opting for BEV segmentation over lane topology for static scene perception, or the training strategies with limited modules’ data.
@@ 5.3 Data Engine
The importance of large-scale and high-quality data for autonomous driving can never be emphasized enough [272]. Establishing a data engine with an automatic labeling pipeline [273] could greatly facilitate the iterative development of both data and models. The data engine for autonomous driving, especially modular end-to-end planning systems, needs to streamline the process of annotating high-quality perception labels with the aid of large perception models in an automatic way. It should also support mining hard/corner cases, scene generation, and editing to facilitate the data-driven evaluations discussed in Sec. 3.2 and promote diversity of data and the generalization ability of models (Sec. 4.9). A data engine would enable autonomous driving models to make consistent improvements.
@@ 5.4 Foundation Model
Recent advancements in foundation models in both language [166, 167, 274] and vision [275, 273, 276] have proved that large-scale data and model capacity can unleash the immense potential of AI in high-level reasoning tasks. The paradigm of finetuning [277] or prompt learning [278], optimization in the form of self-supervised reconstruction [279] or contrastive pairs [165], etc., are all applicable to the end-to-end driving domain. However, we contend that the direct adoption of LLMs for driving might be tricky. The output of an autonomous agent requires steady and accurate measurements, whereas the generative output in language models aims to behave like humans, irrespective of its accuracy. A feasible solution to develop a “foundation” driving model is to train a world model that can forecast the reasonable future of the environment, either in 2D, 3D, or latent space. To perform well on downstream tasks like planning, the objective to be optimized for the model needs to be sophisticated enough, beyond frame-level perception.
6 Conclusion and Outlook
In this survey, we provide an overview of fundamental methodologies and summarize various aspects of simulation and benchmarking. We thoroughly analyze the extensive literature to date, and highlight a wide range of critical challenges and promising resolutions.
Outlook. The industry has dedicated considerable effort over the years to develop advanced modular-based systems capable of achieving autonomous driving on highways. However, these systems face significant challenges when confronted with complex scenarios, e.g., inner-city streets and intersections. Therefore, an increasing number of companies have started exploring end-to-end autonomous driving techniques specifically tailored for these environments. It is envisioned that with extensive high-quality data collection, large-scale model training, and the establishment of reliable benchmarks, the end-to-end approach will soon surpass modular stacks in terms of performance and effectiveness. In summary, end-to-end autonomous driving faces great opportunities and challenges simultaneously, with the ultimate goal of building generalist agents. In this era of emerging technologies, we hope this survey could serve as a starting point to shed new light on this domain.
@@ Acknowledgments
Li Chen, Penghao Wu, and Hongyang Li were partially supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). Andreas Geiger and Bernhard Jaeger were supported by the ERC Starting Grant LEGO-3D (850533), the BMWi in the project KI Delta Learning (project number 19A19013O) and the DFG EXC number 2064/1 - project number 390727645. Kashyap Chitta was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Bernhard Jaeger and Kashyap Chitta.